OpenStudy (anonymous):

simplify 32 ^ (-3/5)

OpenStudy (zehanz):

You must know the following if you are to do these kind of problems:\[a ^{\frac{ b }{ c }}=\sqrt[c]{a^b}=(\sqrt[c]{a})^b \]Also:\[a^{-b}=\frac{ 1 }{ a^b }\]If you look good, you can see that both rules have to be applied here... Can you see what the result is?

OpenStudy (eyust707):

here is how you apply those rules to this problem:

OpenStudy (eyust707):

Notice the second rule refers to the negative exponent. Whenever we see a negative exponent we are allowed to bring it down and make it positive like this: